OpenStudy (anonymous):

Choose the correct simplification 2a^0

OpenStudy (anonymous):

anything raised to 0 is equal to 1 (note: that only affects the base)

OpenStudy (anonymous):

I believe the answer is 1

OpenStudy (johnweldon1993):

Remember anything raised to the 0 power...=1 *just notice....only 'a' is raised to the 0 power*

OpenStudy (anonymous):

So what's the legit answer is it 2a or 1

OpenStudy (johnweldon1993):

Well you have 2a^0 only 'a' is raised to the 0...and anything raised to the 0 is = 1 so you have 2 * 1...which = ?

OpenStudy (anonymous):

1

OpenStudy (johnweldon1993):

2 times 1... = ?

OpenStudy (anonymous):

Equals 1

OpenStudy (anonymous):

Lol sorry 2

OpenStudy (johnweldon1993):

2 times 1 equals 2 Anything times 1...= that number 100 * 1 = 100 5 * 1 = 5 etc...

OpenStudy (anonymous):

So the answer is 2?

OpenStudy (johnweldon1993):

That is correct

OpenStudy (anonymous):

Thanks I have 3 more question I think I know them but I'm not shre

OpenStudy (anonymous):

Choose the correct simplification for x^2/x^5 and the correct simplification for x^6/x^3

OpenStudy (johnweldon1993):

Let me see what you think first and we'll work from there...

OpenStudy (anonymous):

Okay I think that the first one is x^3

OpenStudy (anonymous):

And so is the second one

OpenStudy (johnweldon1993):

Well remember...when dividing exponents with the same base...you are subtracting the exponents and leaving the base alone....so the first one x^2 / x^5 = x^(2 - 5) = x^(?)

OpenStudy (anonymous):

-3

OpenStudy (johnweldon1993):

Right...so the first one is x^-3 The second one.. you have posted above to be x^3...is correct

OpenStudy (anonymous):

Okay but it's a multiple question and -3 is not there

OpenStudy (anonymous):

So is it just x^3 for the first one?

OpenStudy (johnweldon1993):

No...unless you meant x^5/x^2 ?

OpenStudy (anonymous):

Itswriitin the other way around

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

What about (x^9)^2

Directrix (directrix):

Which one of these problems are we doing: Choose the correct simplification for x^2/x^5 OR the correct simplification for x^6/x^3 ?

OpenStudy (anonymous):

The first one and the one I just wrote

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

x^2/x^5 = 1/x^3